Phoenix Company common stock is currently selling $20 per share. Security analysts at Smith Blarney have have assigned the following probability distribution to the price of (and rate of return on) Phoenix stock one year from now:

Price Rate of Return Probability
$16 -20% 0.25
$20 0% 0.30
$24 +20% 0.25
$28 +40% 0.20

1. Assuming that Phoenix is not expected to pay any dividends during the coming year, determine the standard deviation of possible rates of return on Phoenix stock (to the nearest tenth of a percent).

2. Assuming that Phoenix is not expected to pay any dividends during the the coming year, determine the coefficient of variation for the rate of return on Phoenix stock.

**If TI-83 calculator option is available please write out calculator steps**
Answer #1

1. Expected Rate of Return of Phoenix =0.25*-20%+0.30*0%+0.25*20%+0.20*40% =8%
Standard Deviation =(0.25*(-20%-8%)^2+0.30*(0%-8%)^2+0.25*(20%-8%)^2+0.20*(40%-8%)^2)^0.5 =21.35%

2.Coefficient of Variation= Standard Deviation/Expected Return =21.35%/8% =2.669 or 2.67